# Neural Audio Codecs for Prompt-Driven Universal Sound Separation

### v3

@@ 5 Conclusion
Future Work. We will (i) extend separation to higher bandwidths and stereo/spatial audio (e.g., 48 kHz stereo Encodec; HO–DirAC Hold et al. (2024), SpatialCodec Xu et al. (2024)), (ii) broaden prompt coverage with temporal/relational and referring-expression cues, (iii) explore stronger separator objectives, (iv) extend conditioning beyond text to audio/image and mixed prompts via the same FiLM interface.
Appendix A Design Rationale: FiLM-Conditioned Masking in NAC Latent Space
@@
A.1 Why NAC latents (vs. spectrograms) and how the codec prior enables separation?
@@
A.2 Transformer-based masker vs. decoder-style generation (specific design choice).
@@
A.3 Why We Integrated FiLM Conditioning into NAC-Based Separation
@@
A.4 Why report results on $Z$ (continuous latents) and how to extend to code streams $A\!\to\!E$
@@
A.5 Deployment advantages of CodecSep over spectrogram-domain separators (AudioSep)
@@
In the common (code-stream) case, CodecSep removes both decode and re-encode on the server. Moreover, within the separator, CodecSep operates on $Z/E$ with $|Z|\ll|X|$ (cf. Sec. A), reducing activation memory and bandwidth throughout attention and MLP blocks.
@@
  • Eliminates redundant codec cycles in server workflows: With code streams, we reconstruct embeddings $E\approx Z$ by codebook lookup (Sec. A.4), avoiding server-side decode/re-encode; only the edge decodes the final stems.
@@
Appendix B Dataset Details
B.1 Divide and Remaster v 2.0 (dnr-v2)
@@ Appendix B Dataset Details
B.2 Open-Domain Benchmarks
@@
Appendix C Training Details
@@
Since TDANet and CodecFormer were originally developed for speech separation, we re-trained newly initialized versions on the dnr-v2 training set using the same configuration as CodecSep. For AudioSep, we evaluate the publicly available pre-trained model—trained on diverse datasets—and versions re-trained on dnr-v2 and AudioCaps using our setup for consistency. We also include SDCodec, using the official pre-trained models made available by the authors. To ensure a fair comparison, all inputs to TDANet and AudioSep undergo codec processing with a full-band stereo-capable $48$ kHz EnCodec during training and inference. This accounts for codec-induced distortions and artifacts, reflecting realistic deployment scenarios where audio is typically processed through compression pipelines in cloud-based systems.
Appendix D Cross-Benchmark Performance: AudioCaps and dnr-v2.
@@
AudioSep $\mathbf{-2.51^{\pm 12.14}}$ $\mathbf{2.44^{\pm 1.08}}$
AudioSep + dnr-v2 (zero-shot) $6.44^{\pm 11.48}$ $2.33^{\pm 1.08}$
CodecSep + dnr-v2 (zero-shot) $\mathbf{-6.09^{\pm 11.62}}$ $2.24^{\pm 1.16}$
AudioSep + AudioCaps $9.17^{\pm 18.71}$ $2.29^{\pm 1.11}$
CodecSep + AudioCaps $\mathbf{-6.19^{\pm 10.58}}$ $2.14^{\pm 1.00}$
@@
AudioSep + AudioCaps SI-SDR $14.86^{\pm 23.08}$ $7.11^{\pm 25.80}$ $14.63^{\pm 23.26}$
(zero-shot) ViSQOL $2.36^{\pm 0.71}$ $2.43^{\pm 0.70}$ $2.15^{\pm 0.79}$
CodecSep + AudioCaps SI-SDR $8.46^{\pm 2.78}$ $2.47^{\pm 2.91}$ $5.93^{\pm 4.33}$
(zero-shot) ViSQOL $2.27^{\pm 0.53}$ $2.56^{\pm 0.47}$ $2.06^{\pm 0.72}$
Appendix E Discussion of Relative–Gain Summaries.
@@
Appendix F Bandwidth scaling: extending CodecSep to 48 kHz full-band
@@
AudioSep $32$ kHz SI-SDR $2.46^{\pm 4.06}$ $4.92^{\pm 4.21}$ $0.34^{\pm 5.39}$
(zero-shot) ViSQOL $2.86^{\pm 0.63}$ $3.11^{\pm 0.56}$ $\mathbf{2.63^{\pm 0.77}}$
AudioSep + dnr-v2 $32$ kHz SI-SDR $5.55^{\pm 2.89}$ $7.68^{\pm 3.0}$ $4.66^{\pm 3.68}$
ViSQOL $2.59^{\pm 0.57}$ $2.49^{\pm 0.37}$ $2.32^{\pm 0.7}$
CodecSep + dnr-v2 $16$ kHz SI-SDR $\mathbf{1.15^{\pm 3.29}}$ $\mathbf{9.97^{\pm 2.92}}$ $\mathbf{0.89^{\pm 4.22}}$
(DAC Backbone) ViSQOL $\mathbf{2.86^{\pm 0.57}}$ $\mathbf{3.14^{\pm 0.45}}$ $2.33^{\pm 0.73}$
CodecSep + dnr-v2 $48$ kHz SI-SDR $2.8^{\pm 3.5}$ $5.36^{\pm 2.36}$ $0.46^{\pm 3.83}$
(EnCodec Backbone) ViSQOL $2.39^{\pm 0.5}$ $2.63^{\pm 0.41}$ $2.37^{\pm 0.65}$
Appendix G Further Studies: Reconstruction Performance.
@@
TDANet SI-SDR $3.26^{\pm 7.78}$ $\mathbf{7.96^{\pm 5.29}}$ $\mathbf{11.07^{\pm 3.32}}$ $\mathbf{4.70^{\pm 5.11}}$
ViSQOL $3.93^{\pm 0.35}$ $\mathbf{4.20^{\pm 0.46}}$ $\mathbf{4.51^{\pm 0.32}}$ $\mathbf{4.09^{\pm 0.39}}$
CodecFormer SI-SDR $47.55^{\pm 9.51}$ $47.11^{\pm 10.97}$ $47.77^{\pm 9.65}$ $48.21^{\pm 9.87}$
ViSQOL $1.02^{\pm 0.07}$ $4^{\pm 0.12} 1.01±0$$^{\pm 0.06} 1.1$$16^{\pm 0.47}$
CodecSep + dnr-v2 SI-SDR $3.42^{\pm 1.85}$ $4.12^{\pm 3.97}$ $6.15^{\pm 2.87}$ $0.83^{\pm 5.16}$
(unguided, 3-stem) ViSQOL $3.23^{\pm 0.20}$ $3.02^{\pm 0.33}$ $3.47^{\pm 0.24}$ $3.23^{\pm 0.46}$
SDCodec SI-SDR $\mathbf{6.98^{\pm 2.49}}$ $7.65^{\pm 4.60}$ $8.28^{\pm 3.26}$ $2.54^{\pm 5.65}$
ViSQOL $\mathbf{4.29^{\pm 0.15}}$ $03^{\pm 0.28} 4.4$$44^{\pm 0.15}$ $3.98^{\pm 0.34}$
@@
AudioSep SI-SDR $5.53^{\pm 1.96}$ $4.69^{\pm 5.36}$ $10.97^{\pm 2.99}$ $1.99^{\pm 5.68}$
(zero-shot) ViSQOL $4.06^{\pm 0.38}$ $3.75^{\pm 0.65}$ $4.57^{\pm 0.13}$ $3.19^{\pm 0.77}$
AudioSep + dnr-v2 SI-SDR $6.47^{\pm 2.26}$ $7.99^{\pm 4.55}$ $8.13^{\pm 3.35}$ $2.29^{\pm 5.95}$
ViSQOL $4.20^{\pm 0.18}$ $4.12^{\pm 0.21}$ $3.03^{\pm 0.29}$ $3.81^{\pm 0.47}$
CodecSep + dnr-v2 SI-SDR $4.06^{\pm 2.06}$ $3.91^{\pm 3.93}$ $6.10^{\pm 2.86}$ $0.67^{\pm 5.29}$
ViSQOL $74^{\pm 0.22} 3.3$$37^{\pm 0.33}$ $3.83^{\pm 0.24}$ $3.54^{\pm 0.44}$
CodecSep + dnr-v2 SI-SDR $\mathbf{12.24^{\pm 2.42}}$ $\mathbf{12.58^{\pm 3.81}}$ $\mathbf{13.59^{\pm 2.59}}$ $\mathbf{8.67^{\pm 4.17}}$
(ablate Masker) ViSQOL $\mathbf{4.44^{\pm 0.14}}$ $\mathbf{4.13^{\pm 0.31}}$ $\mathbf{3.85^{\pm 0.34}}$ $3.76^{\pm 0.54}$
AudioSep + AudioCaps SI-SDR $6.67^{\pm 2.52}$ $8.10^{\pm 4.63}$ $8.39^{\pm 3.21}$ $2.42^{\pm 6.12}$
(zero-shot) ViSQOL $4.23^{\pm 0.19}$ $4.12^{\pm 0.21}$ $4.17^{\pm 0.21}$ $\mathbf{3.82^{\pm 0.46}}$
CodecSep + AudioCaps SI-SDR $0.59^{\pm 1.89}$ $0.24^{\pm 5.15}$ $11.03^{\pm 5.21}$ $1.23^{\pm 4.84}$
(zero-shot) ViSQOL $3.33^{\pm 0.23}$ $2.91^{\pm 0.64}$ $1.69^{\pm 0.48}$ $3.44^{\pm 0.41}$

### v4

@@ 5 Conclusion
Future Work. We plan to (i) extend separation to higher bandwidths and stereo/spatial audio (e.g., 48 kHz stereo Encodec; HO–DirAC Hold et al. (2024), SpatialCodec Xu et al. (2024)); (ii) broaden prompt coverage to include temporal, relational, and referring-expression cues; (iii) explore stronger separator objectives; (iv) extend conditioning beyond text to audio, image, and mixed prompts through the same FiLM interface; and (v) pursue interpretability analyses. In particular, the predicted masks and FiLM responses indicate how CodecSep selectively activates latent components, suggesting a clear direction for future work involving mask visualization and reconstruction-consistency checks.
Appendix A Failure Modes of PIT and MixIT for Universal Sound Separation
Permutation-Invariant Training (PIT) and Mixture-Invariant Training (MixIT) have historically been effective for closed-domain separation tasks where the number of underlying sources is known, fixed, or varies within a narrow, well-defined range. However, their underlying assumptions lead to structural limitations when applied to open-domain universal source separation (USS), where mixtures may contain an arbitrary and potentially large number of heterogeneous sound events. In this section, we summarize the key failure modes observed when training with PIT/MixIT to extend fixed-stem models to open-domain mixtures.
A core limitation of PIT/MixIT is the requirement to specify a maximum number of output sources, denoted by $N$. During training, the model produces exactly $N$ outputs for every mixture, and the PIT or MixIT objective establishes a correspondence between these outputs and the underlying reference sources (or intermediate MixIT partitions). This design is brittle in scenarios where the true number of sources varies widely. When the mixture contains more than $N$ sources—a frequent occurrence in open-domain audio—the model has no mechanism to create additional outputs. Instead, it suffers source collision and collapses multiple sources into a single output stem, resulting in unavoidable leakage, loss of fine structure, and a sharp degradation in separation quality. The post-hoc identification step cannot recover the missing sources, because the model never produced separate representations for them in the first place; those sources simply do not exist within the model’s output space.
Conversely, when the mixture contains fewer than $N$ sources, the model is still obligated to return $N$ outputs. This mismatch introduces new problems: several outputs correspond to no actual source and become “inactive” stems, while others may capture residual background energy or hallucinated content. These false positives degrade metrics such as SI-SDR and create ambiguity during evaluation because the model does not encode which stems are meant to be meaningful. Such outputs also make deployment difficult, as downstream systems must decide which stems to trust and which to ignore.
The reliance on a fixed maximum number of sources $N$ also places a heavy burden on both training stability and computational cost. As $N$ increases, the permutation space in PIT expands combinatorially, and MixIT assignments become increasingly complex, making training slow, unstable, and in many cases prone to divergence. In open-domain datasets such as dnr-v2, mixtures may contain eight or more concurrent sound events, forcing PIT/MixIT baselines to adopt impractically large values of $N$ to avoid source collisions. In practice, such configurations are computationally prohibitive and empirically unreliable.
These limitations collectively illustrate why PIT and MixIT, despite their historical success in speech separation and other closed-set tasks, are poorly suited for open-domain universal separation. Their fixed-output architecture is fundamentally mismatched to real-world mixtures that contain highly variable and unpredictable numbers of sources. In contrast, CodecSep bypasses this bottleneck entirely through free-form text-guided inference: the model extracts only the requested source category, emits no unused stems, and scales naturally to mixtures with arbitrary levels of overlap. This flexibility enables CodecSep to support both closed-set and open-domain use cases, while also providing a foundation for future extensions to fine-grained extraction of individual speaker stems, instrument stems, or sound-effect stems.
Appendix B Design Rationale: FiLM-Conditioned Masking in NAC Latent Space
@@
B.1 Why NAC latents (vs. spectrograms) and how the codec prior enables separation?
@@
B.2 Transformer-based masker vs. decoder-style generation (specific design choice).
@@
B.3 Why We Integrated FiLM Conditioning into NAC-Based Separation
@@
B.4 Why report results on $Z$ (continuous latents) and how to extend to code streams $A\!\to\!E$
@@
B.5 Deployment advantages of CodecSep over spectrogram-domain separators (AudioSep)
@@
In the common (code-stream) case, CodecSep removes both decode and re-encode on the server. Moreover, within the separator, CodecSep operates on $Z/E$ with $|Z|\ll|X|$ (cf. Sec. B), reducing activation memory and bandwidth throughout attention and MLP blocks.
@@
  • Eliminates redundant codec cycles in server workflows: With code streams, we reconstruct embeddings $E\approx Z$ by codebook lookup (Sec. B.4), avoiding server-side decode/re-encode; only the edge decodes the final stems.
@@
Appendix C Dataset Details
C.1 Divide and Remaster v 2.0 (dnr-v2)
@@ Appendix C Dataset Details
C.2 Open-Domain Benchmarks
@@
Appendix D Evaluation Details
Figure 3: Evaluation workflow for dnr-v2. Each mixture contains multi-source stems: speech (often multi-speaker), music (multi-instrument), and SFX ($\geq 3$ overlapping events). Fixed-stem baselines predict a fixed set of outputs (e.g., 3 stems), whereas CodecSep and other text-guided models generate only the prompted source. Speech and music are evaluated using generic prompts, while SFX uses long-form compositional prompts listing all SFX events in each mixture. Extracted signals are compared with ground-truth category stems using SI-SDR and ViSQOL.
D.1 Divide and Remaster v 2.0 (dnr-v2)
The dnr-v2 benchmark presents a challenging open-domain separation setting: although the dataset provides three category labels—speech, music, and sound effects—each category represents a multi-source stem. A single mixture frequently contains five to ten underlying acoustic sources, including overlapping speakers, multiple musical instruments, and several sound-effect events occurring either concurrently or in sequence. The three reported stems are therefore semantic groupings that support interpretability and reproducibility, rather than an indication that the mixture contains only three sources. Any evaluation methodology must respect this structure. Figure 3 provides an overview of our dnr-v2 evaluation workflow and highlights how these multi-source stems are handled across fixed-stem un-guided and text-guided models.
For fixed-stem unguided architectures, evaluation is performed by mapping each predicted output stem to one of the three ground-truth stems and computing SI-SDR and ViSQOL on a per-category basis. Importantly, we do not employ PIT or MixIT training objectives for these baselines; instead, we train dedicated three-stem models that directly predict speech, music, and SFX stems.
Text-guided models, including CodecSep, follow a fundamentally different inference and evaluation paradigm. Speech and music stems are recovered using generic prompts (“speech,” “music”), which reliably capture their multi-source content. In contrast, SFX stems require mixture-specific prompts because sound effects span a wide and open-domain label space. For each mixture, we use a long-form compositional prompt enumerating all SFX events present in the ground truth. This ensures that the model has sufficient semantic context to extract the full SFX stem. The number of sfx events present in a mixture varies considerably. We additionally perform an ambiguous-prompt evaluation, where deliberately underspecified prompts for speech and music are used to assess robustness to vague or incomplete semantic queries. After inference, the extracted waveform for each category is directly compared with the corresponding ground-truth stem using SI-SDR and ViSQOL. This evaluation design ensures fairness between fixed-stem and text-guided systems while faithfully reflecting the multi-source structure of dnr-v2 mixtures
Figure 4: Evaluation workflow for the standardized three-source benchmarks (AudioCaps, ESC-50, Clotho, VGGSound, and AudioSet-eval). Following prior USS protocols, each mixture is constructed by combining three isolated events drawn from distinct classes. For each class, the corresponding textual prompt is supplied to the separator (e.g., “dog barking,” “gun shot,” “motor vehicle”), and the extracted signal is compared with the ground-truth isolated source using SI-SDR and ViSQOL.
D.2 Open-Domain Benchmarks
For AudioCaps, ESC-50, Clotho, VGGSound, and AudioSet-eval, we adopt the standardized three-source mixture protocol. Following established practice, each mixture is constructed by combining three isolated events drawn from different classes. Each source is then extracted by the text-guided models using its corresponding textual caption as prompt, and evaluation metrics are computed against the ground-truth isolated audio. Figure 4 illustrates the evaluation workflow used for these benchmarks, highlighting how class-specific prompts are applied and how the resulting predictions are matched against the ground-truth isolated sources. Although these datasets do not reflect the complex multi-source structure of dnr-v2, the standardized 3-way protocol enables direct benchmarking against prior work (AudioSep) under consistent conditions.
Appendix E Training Details
@@
Since TDANet and CodecFormer were originally designed for speech separation, we re-train newly initialized 3-stem versions on the dnr-v2 training set using the same configuration as CodecSep. We also train a 3-stem Sudo rm-rf model to compare against compute-efficient separators. For AudioSep, we evaluate both the publicly available pretrained model—trained on diverse datasets—and versions re-trained on dnr-v2 and AudioCaps for consistency. We include SDCodec using the official pretrained checkpoints released by the authors. Finally, we incorporate the USS-pretrained variant of BiModalSS and re-train SudoRmRf+FiLM on dnr-v2 with CLAP text conditioning. To ensure a fair comparison, all inputs to TDANet, Sudo rm-rf, AudioSep, BiModalSS, and Sudo rm-rf+FiLM undergo codec processing with a full-band stereo-capable $48$ kHz EnCodec during training and inference. This accounts for codec-induced distortions and artifacts, reflecting realistic deployment scenarios where audio is typically processed through compression pipelines in cloud-based systems.
Appendix F Robustness to prompt paraphrasing.
To probe lexical sensitivity, we re-evaluated both CodecSep + dnr-v2 and AudioSep + dnr-v2 on the dnr-v2 test split by replacing the generic training-time prompts for speech and music with three unseen paraphrases per class—speech: {“spoken voice”, “human conversation”, “people talking”}; music: {“instrumental music”, “band playing”, “melody with instruments”}. This constitutes a zero-shot paraphrase generalization test: models are trained with generic category cues but must respond to synonymic, potentially broader descriptors at inference. We observe a consistent qualitative pattern (cf. Table 6): (i) both systems exhibit the expected degradation when moving from generic to paraphrased prompts, confirming that lexical ambiguity weakens query–audio alignment; (ii) CodecSep degrades more gracefully overall, maintaining stronger separation and perceptual quality for speech, and retaining a small but persistent advantage for music; and (iii) the gap between models narrows under paraphrasing, yet the relative ranking is preserved, suggesting that FiLM-conditioned masking over structured codec latents confers robustness to synonym-level shifts. Notably, this study isolates lexical paraphrases; we did not incorporate paraphrases with explicit temporal qualifiers (e.g., “applause follows a song”), which we leave to future work.
Model Metric ($\uparrow$) Music Speech
AudioSep + dnr-v2 SI-SDR $6.4^{\pm 3.29}$ $4.1^{\pm 3.77}$
ViSQOL $2.5^{\pm 0.57}$ $2.6^{\pm 0.47}$
CodecSep + dnr-v2 SI-SDR $\mathbf{-5.6^{\pm 3.61}}$ $\mathbf{4.2^{\pm 4.18}}$
ViSQOL $\mathbf{2.6^{\pm 0.58}}$ $\mathbf{2.7^{\pm 0.51}}$
Appendix G Cross-Benchmark Performance: AudioCaps and dnr-v2.
@@
AudioSep $\mathbf{-2.5^{\pm 12.14}}$ $\mathbf{2.4^{\pm 1.08}}$
AudioSep + dnr-v2 (zero-shot) $6.4^{\pm 11.48}$ $2.3^{\pm 1.08}$
CodecSep + dnr-v2 (zero-shot) $\mathbf{-6.1^{\pm 11.62}}$ $2.2^{\pm 1.16}$
AudioSep + AudioCaps $9.2^{\pm 18.71}$ $2.3^{\pm 1.11}$
CodecSep + AudioCaps $\mathbf{-6.2^{\pm 10.58}}$ $2.1^{\pm 1.00}$
@@
AudioSep + AudioCaps SI-SDR $14.9^{\pm 23.08}$ $7.1^{\pm 25.80}$ $14.6^{\pm 23.26}$
(zero-shot) ViSQOL $2.4^{\pm 0.71}$ $4^{\pm 0.70} 2.2$$2^{\pm 0.79}$
CodecSep + AudioCaps SI-SDR $8.5^{\pm 2.78}$ $2.5^{\pm 2.91}$ $5.9^{\pm 4.33}$
(zero-shot) ViSQOL $2.3^{\pm 0.53}$ $2.6^{\pm 0.47}$ $2.1^{\pm 0.72}$
Appendix H Discussion of Relative–Gain Summaries.
@@
Appendix I Bandwidth scaling: extending CodecSep to full-band audio
@@
AudioSep $32$ kHz SI-SDR $2.5^{\pm 4.06}$ $4.9^{\pm 4.21}$ $0.3^{\pm 5.39}$
(zero-shot) ViSQOL $2.9^{\pm 0.63}$ $3.1^{\pm 0.56}$ $\mathbf{2.6^{\pm 0.77}}$
AudioSep + dnr-v2 $32$ kHz SI-SDR $5.6^{\pm 2.89}$ $7.7^{\pm 3.0}$ $4.5^{\pm 3.68}$
ViSQOL $2.6^{\pm 0.57}$ $2.5^{\pm 0.37}$ $2.3^{\pm 0.7}$
CodecSep + dnr-v2 $16$ kHz SI-SDR $\mathbf{1.2^{\pm 3.29}}$ $\mathbf{10.0^{\pm 2.92}}$ $\mathbf{0.9^{\pm 4.22}}$
(DAC Backbone) ViSQOL $\mathbf{2.9^{\pm 0.57}}$ $\mathbf{3.1^{\pm 0.45}}$ $2.3^{\pm 0.73}$
CodecSep + dnr-v2 $24$ kHz SI-SDR $0.2^{\pm 3.3}$ $8.8^{\pm 2.9}$ $0.6^{\pm 4.2}$
(DAC Backbone) ViSQOL $2.7^{\pm 0.56}$ $3.0^{\pm 0.44}$ $2.3^{\pm 0.72}$
CodecSep + dnr-v2 $44.1$ kHz SI-SDR $2.3^{\pm 3.27}$ $5.9^{\pm 2.48}$ $0.3^{\pm 3.79}$
(DAC Backbone) ViSQOL $2.5^{\pm 0.46}$ $2.7^{\pm 0.42}$ $2.4^{\pm 0.68}$
CodecSep + dnr-v2 $48$ kHz SI-SDR $2.8^{\pm 3.5}$ $5.4^{\pm 2.36}$ $0.5^{\pm 3.83}$
(EnCodec Backbone) ViSQOL $2.4^{\pm 0.5}$ $2.6^{\pm 0.41}$ $2.4^{\pm 0.65}$
Appendix J Further Studies: Reconstruction Performance.
@@
TDANet SI-SDR $3.3^{\pm 7.78}$ $\mathbf{8.0^{\pm 5.29}}$ $\mathbf{11.1^{\pm 3.32}}$ $\mathbf{4.7^{\pm 5.11}}$
ViSQOL $3.9^{\pm 0.35}$ $\mathbf{4.2^{\pm 0.46}}$ $\mathbf{4.5^{\pm 0.32}}$ $\mathbf{4.1^{\pm 0.39}}$
CodecFormer SI-SDR $47.6^{\pm 9.51}$ $47.1^{\pm 10.97}$ $47.8^{\pm 9.65}$ $48.2^{\pm 9.87}$
ViSQOL $^{\pm 0.07} 1.0±0.1$$^{\pm 0.12}$ $1.0^{\pm 0.06}$ $1.2^{\pm 0.47}$
CodecSep + dnr-v2 SI-SDR $3.4^{\pm 1.85}$ $4.1^{\pm 3.97}$ $6.2^{\pm 2.87}$ $0.8^{\pm 5.16}$
(unguided, 3-stem) ViSQOL $3.2^{\pm 0.20}$ $3.0^{\pm 0.33}$ $3.5^{\pm 0.24}$ $3.2^{\pm 0.46}$
SDCodec SI-SDR $\mathbf{7.0^{\pm 2.49}}$ $7.7^{\pm 4.60}$ $8.3^{\pm 3.26}$ $2.5^{\pm 5.65}$
ViSQOL $\mathbf{4.3^{\pm 0.15}}$ $0^{\pm 0.28} 4.4$$4^{\pm 0.15}$ $4.0^{\pm 0.34}$
@@
AudioSep SI-SDR $5.5^{\pm 1.96}$ $4.7^{\pm 5.36}$ $11.0^{\pm 2.99}$ $2.0^{\pm 5.68}$
(zero-shot) ViSQOL $4.1^{\pm 0.38}$ $3.8^{\pm 0.65}$ $4.6^{\pm 0.13}$ $3.2^{\pm 0.77}$
AudioSep + dnr-v2 SI-SDR $6.5^{\pm 2.26}$ $8.0^{\pm 4.55}$ $8.1^{\pm 3.35}$ $2.3^{\pm 5.95}$
ViSQOL $4.2^{\pm 0.18}$ $4.1^{\pm 0.21}$ $3.0^{\pm 0.29}$ $3.8^{\pm 0.47}$
CodecSep + dnr-v2 SI-SDR $4.1^{\pm 2.06}$ $3.9^{\pm 3.93}$ $6.1^{\pm 2.86}$ $0.7^{\pm 5.29}$
ViSQOL $3.7^{\pm 0.22}$ $3.4^{\pm 0.33}$ $3.8^{\pm 0.24}$ $3.5^{\pm 0.44}$
CodecSep + dnr-v2 SI-SDR $\mathbf{12.2^{\pm 2.42}}$ $\mathbf{12.6^{\pm 3.81}}$ $\mathbf{13.6^{\pm 2.59}}$ $\mathbf{8.7^{\pm 4.17}}$
(ablate Masker) ViSQOL $\mathbf{4.4^{\pm 0.14}}$ $\mathbf{4.1^{\pm 0.31}}$ $\mathbf{3.9^{\pm 0.34}}$ $3.8^{\pm 0.54}$
AudioSep + AudioCaps SI-SDR $6.7^{\pm 2.52}$ $8.1^{\pm 4.63}$ $8.4^{\pm 3.21}$ $2.4^{\pm 6.12}$
(zero-shot) ViSQOL $4.2^{\pm 0.19}$ $4.1^{\pm 0.21}$ $4.2^{\pm 0.21}$ $\mathbf{3.8^{\pm 0.46}}$
CodecSep + AudioCaps SI-SDR $0.6^{\pm 1.89}$ $0.2^{\pm 5.15}$ $11.^{\pm 5.21}$ $1.2^{\pm 4.84}$
(zero-shot) ViSQOL $3.3^{\pm 0.23}$ $2.9^{\pm 0.64}$ $1.7^{\pm 0.48}$ $3.4^{\pm 0.41}$